OpenStudy (anonymous):

find the volume of the given solid over the indicated region of the integration. f(x,y) = 2x+3y+2; -2<=x<=4; 2<=y<=5

OpenStudy (anonymous):

This is how I've started \[\int\limits_{0}^{4} \int\limits_{2}^{5}2x +3y +2dydx -\int\limits_{-2}^{0}\int\limits_{2}^{5}2x +3y +2dydx\]

OpenStudy (anonymous):

Solved for the inside integral of both first \[\int\limits_{2}^{5}(2x +3y +2) dy = 2xy + \frac{ 3 }{ 2 }y^2 +2y |\] \[ y=5: 10x+\frac{ 75 }{ 2 }+10\] \[-(y=2:4x+48+4)\] \[= 6x- 9/2\]

OpenStudy (anonymous):

Am I on the right track?

OpenStudy (anonymous):

why take a negative sign? it can be -2 to 4. \[V=\int\limits_{-2}^{4}\left( 6x-\frac{ 9 }{ 2 } \right)dx\]

OpenStudy (anonymous):

Finding volume. Wouldn't the negative part be subtracted from the total that way?

OpenStudy (anonymous):

i don't think so.

OpenStudy (anonymous):

\[V=\left[ \frac{ 6 x^2 }{ 2 }-\frac{ 9 }{ 2 }x \right]~from~-2~\to~4=3\left( 4^2-(-2)^2 \right)-\frac{ 9 }{ 2}\left( 4-(-2) \right)\]

OpenStudy (anonymous):

wolfram alpha says 261, this says 9, I found 43 my way (which was wrong)

ganeshie8 (ganeshie8):

your splitting doesn't look correct

ganeshie8 (ganeshie8):

the solid is living both above and below xy plane and you're spluttung because you want to integrate the absolute value, right ?

OpenStudy (anonymous):

right

ganeshie8 (ganeshie8):

can u give me wolfram link

ganeshie8 (ganeshie8):

wolfram is giving you signed volume

ganeshie8 (ganeshie8):

you should get same or more than that if u take absolute value

OpenStudy (anonymous):

Same result, then?

ganeshie8 (ganeshie8):

Yes because the solid is completely above the xy plane in the given region

ganeshie8 (ganeshie8):

no splitting is needed

OpenStudy (anonymous):

Do you know the correct way to set this up?

ganeshie8 (ganeshie8):

`f(x,y) = 2x+3y+2; -2<=x<=4; 2<=y<=5` first notice that \[\large f(x,y)= 2x+3y+2 \ge 0 \\~\\ \text{ when } -2\le x\le 4 \text{ and } 2\le y \le 5 \]

ganeshie8 (ganeshie8):

So \[\int\limits_{-2}^{4} \int\limits_{2}^{5}\Bigg|2x +3y +2\Bigg|dydx =\int\limits_{-2}^{4} \int\limits_{2}^{5}2x +3y +2~dydx \]

ganeshie8 (ganeshie8):

http://gyazo.com/8b19ba6aa78f053b44006987b0a9c7bd

OpenStudy (anonymous):

leads to this?\[\int\limits_{-2}^{4} \left( 2xy +\frac{ 3 }{ 2 } y^2 + 2y from 2\rightarrow5\right)dx\]

ganeshie8 (ganeshie8):

looks good

OpenStudy (anonymous):

\[\int\limits_{-2}^{4}6x-\frac{ 9 }{ 2 }dx\]

ganeshie8 (ganeshie8):

im getting 6x + 75/2

OpenStudy (anonymous):

\[\int\limits_{-2}^{4}6x+\frac{ 75 }{ 2 }dx\] \[\left( 3x^2 +\frac{ 75 }{ 2 }x \right)\] from -2 to 4

ganeshie8 (ganeshie8):

Yep

OpenStudy (anonymous):

=261

OpenStudy (anonymous):

!! Thank you again ganeshie!

ganeshie8 (ganeshie8):

np :)